OpenStudy (anonymous):

If a line with a slope of 4 crosses the y-axis at (0, 1), what is the equation of the line?

OpenStudy (anonymous):

A. y − 4x = 1 B. y + 4x = 1 C. x − 4y = 1 D. x + 4y = 1

OpenStudy (anonymous):

@Preetha

OpenStudy (mertsj):

It it crosses the y axis at (0,1) that means that the y intercept is 1

OpenStudy (anonymous):

Is it a or b?

OpenStudy (mertsj):

Put them both in y = mx+b form and see which one has slope 4 and y intercept 1

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

yes